Session: Feb 2017 (session 1)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Angel Vicente at Sun Mar 12 07:17:46 2017) Minor failures:
Problem recording scan12

Effelsberg (Uwe at Mon Mar 13 09:50:58 2017) Success:
No problems, beside the broken LCP channel.

Onsala (Jun Yang at Wed Mar 15 13:30:54 2017) Success:
No known problems.